Brand X sells 21 oz. bags of mixed nuts contain 29% peanutes. To make their product they combine Brand A mixed nuts which contain 35% peanuts and Brand B mixed nuts which contain 25% peanuts. How much of each do they need to use?
Answer by jorel1380(3719)

21x.29=6.09
m+n=21
.35m+.25n=6.09
35m+25m=609
25m+25n=525
10m=84
m=8.4
n=12.6
You need 8.4 ozs. of 35% peanuts with 12.6 ozs. of 25% peanuts for your mix..